what are vectors and addition and subtraction of vectors

Grade:11

4 Answers

sonika p
6 years ago
vectors are quantities that have magnitude as well as direction. for ex-displacement. when a displacement is defined we say its magnitude may be 5m or so and its direction say, in the west. vectors gives us that precise information about both directions and magnitude. 
addition or subtraction of vectors are simply performing those operations on the components of the vectors. 
for ex- if you first walk 3m towards the west direction and then turn around and walk 2m backwards(east), your effective walk would be just 1m from the initial point and not 5m from the point where you have started walking. 
Hari Sundar Pyda
6 years ago
Vector quantities are quantities which have both magnitude as well as direction. Ex- force, displacement, velocity, etc..Vectors, when represented in three dimensional system, has unit vectors.i^(i inside the cap) for X-axis. j^ for Y-axis.k^ for Z-axis.Ex- F=4i^+6j^+12k^Vectors can be added-They can be shown in graphical representation. They can be added or subtracted using triangular law of vector addition or parallelogram law of vector addition. OrIf A and B are two vectors, their resultant vector= R = root of(A2 + B2 + 2AB cos θ).Vectors cannotbe directly subtracted. Subtraction of a vector B from a vector A is defined as the addition of vector -B (negative of vector B) to vector AThus, A – B = A + (-B)
Harshit Panwar
6 years ago
vectors are the quantites having both magitude and direction.
For eg. acceleration,displacement,force,etc.
Vectors are added by parellogram law of vector addition. In this addition two vectors are arranged in ordered pairs then a line should be drawn form tail of a vector to head to the second vector. The length of diagonal is the addtion of the vectors.
R2=A2+B2  {here ‘R’ is the resultant}
Vector substraction is nothing but vectorA+(-vectorB)
in this method we join vectors at their heads and then by joining their tails by another line gives us the answer.
Gurmeet
6 years ago
There are two different types of quantity :vector and scalar vectors are the quantites having both magitude and direction.For eg. acceleration,displacement,force,etc.Vectors are added by parellogram law of vector addition. In this addition two vectors are arranged in ordered pairs then a line should be drawn form tail of a vector to head to the second vector. The length of diagonal is the addtion of the vectors.R2=A2+B2 {here ‘R’ is the resultant}Vector substraction is nothing but vectorA+(-vectorB)in this method we join vectors at their heads and then by joining their tails by another line gives us the answer.